OpenStudy (anonymous):

Medal and Fan for someone who can check my answer! What is the sum of the geometric sequence 8, -16, 32 . if there are 15 terms?

OpenStudy (anonymous):

First term=a1=8..............r=-16/2=-2.......... ..........a1(1-r^n) S(n)=*********** ...............1-r ............8(1-(-2)^(15)) S(15)=**************** ...................1-(-2) ............8(1-(-2)^(15)) S(15)=**************** ...................1+2 ............8(1-(-2)^(15)) S(15)=**************** ......................3.................. time S(15)=87384..........ANSWER

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

Okay, great. Thank you so much!